# Leveraging Multi-source and Heterogeneous Signals for Fatigue Detection

### v2

@@ Abstract
Fatigue detection plays a critical role in safety-critical applications such as aviation, mining, and long-haul transport. However, most existing methods rely on high-end sensors and controlled environments, limiting their applicability in real-world settings. This paper formally defines a practical yet underexplored problem setting for real-world fatigue detection, where systems operating with context-appropriate sensors aim to leverage knowledge from differently instrumented sources—including those using impractical sensors deployed in controlled environments. To tackle this challenge, we propose a heterogeneous and multi-source fatigue detection framework that adaptively utilizes the available modalities in the target domain while benefiting from the diverse configurations present in source domains. Our experiments, conducted using a realistic field-deployed sensor setup and two publicly available datasets, demonstrate the practicality, robustness, and improved generalization of our approach, paving the practical way for effective fatigue monitoring in sensor-constrained scenarios.
Note to Practitioners—The proposed work is motivated by the growing need for practical and reliable fatigue detection in industrial applications where operator fatigue poses serious safety and productivity risks. While high-performing physiological sensors like EEG provide high-fidelity signals, their use in these environments is limited by practical constraints such as cost, setup complexity, and environment sensitivity. Thus, industry deployment often relies on context-appropriate sensors that may not fully capture fatigue-related states. To resolve this conflict, we propose a multi-source heterogeneous fatigue detection framework that enables field-deployed systems to take advantage of high-fidelity data collected from controlled environments, even if those modalities are unavailable during deployment. Our experiments validate the practicality, robustness, and generalization of our approach, offering a scalable solution to effective fatigue monitoring in sensor-constrained industry settings.
Index Terms:
Fatigue Detection, Sensor Imputation, Data Synthesis
@@ I Introduction
Physiological responses and behavioral patterns of individuals under fatigue provide rich information to understand the impacts of jeopardized human capacity, from which proper mitigation control can be determined. However, the complexity and vulnerability of data acquisition presents several limitations. While the latest advancements of sensor technology have elevated the research of objective fatigue detection to unprecedented levels, concerns about sensor reliability and accessibility remain [1].
@@
Unreliable sensors can result in loss of data, compromising detection accuracy, or, in the worst case, complete malfunction. To address this issue, recent efforts have explored using synthetic data generation strategies to augment limited or incomplete datasets. For instance, generative AI [2] has emerged as a promising solution. Bird et al. (2021) demonstrated the capability of GPT-2 in generating synthetic Electroencephalography (EEG) and Electromyography (EMG), which, when combined with real training data, improved classification accuracy [3]. Similarly, research in [4] employed generative adversarial networks (GAN) to enhance data for motor imagery classification in brain-computer interface systems. Along this direct, [5] proposed Generative-DANN (GDANN), a hybrid model that integrates GANs with a Domain-Adversarial Neural Network (DANN) architecture to address cross-subject EEG variability in fatigue detection. By aligning feature distributions across subjects, GDANN improved generalizability and achieved high accuracy in cross-subject fatigue classification tasks. When generating synthetic data for fatigue detection, the process must consider inherent correlations of physiological modalities to collaboratively determine fatigue. To that end, RAINDROP [6] provided a promising approach by using graph neural networks to model these correlations in irregularly sampled and multivariate time-series data.
@@
Although the aforementioned approaches address the challenge of sensor unreliability through data augmentation, few efforts have tackled the equally critical issue of sensor accessibility. In many real-world scenarios, high-end sensors, while reliable and accurate, remain prohibitively expensive or impractical for broad deployment. As a result, studies involving these sensors are often confined to simulated and controlled environments. For example, high-resolution EEG was deployed in a simulated environment, together with heart rate and eye blinks rate, to analyze mental fatigue or drowsiness during car driving [7]. However, such simulation-based setups often fail to capture the complexity and uncertainty of actual driving conditions [8]. Moreover, these sensors—such as EEG, imaging systems, and pupil dilation monitors—are highly sensitive to environmental factors, including noise, lighting, and field-of-view constraints. These sensitivities complicate data collection and analysis in real-world scenarios, ultimately degrading model accuracy and robustness [9]. This gap presents a compelling opportunity: Can we transfer the high-fidelity knowledge extracted from controlled environments—where certain sensors are deployed but unsuitable for real-world use—to practical settings that rely on a different set of context-appropriate sensors? Instead of attempting to replicate the sensor configurations used in laboratory settings, we hypothesize that a novel domain-adaptive framework can be developed to align heterogeneous signal representations, enabling real-world fatigue detection systems to learn from and be enhanced by control-environment data, even if they originate from entirely different sources. Guided by this bold idea, this paper presents a practical fatigue detection solution, and makes the following contributions:
@@
  • A practical yet underexplored problem setting for real-world fatigue detection is formally defined. This definition reflects the need to operate fatigue detection systems with context-appropriate sensors in the field and presents the opportunity to augment these systems with knowledge transferred from differently instrumented sensor sources. It ultimately lays the foundation for developing models that are robust to modality heterogeneity, incompleteness, and domain shift.
@@
  • A heterogeneous and multi-source fatigue detection framework is proposed as a solution to this problem, supported by a thorough theoretical analysis. In particular, we explore the conditions under which knowledge from differently instrumented source domains can be effectively leveraged to support fatigue prediction in a target domain with incomplete and mismatched sensor modalities. Building on these insights, we then develop a model that adaptively leverages the available modalities in the target domain while benefiting from the diverse modality configurations present across multiple source domains. The design enables accurate fatigue prediction in practical settings, even under constrained sensor configurations, resulting in improved generalization and overall performance.
  • The proposed framework is validated through a practical sensor setup together with two publicly available datasets, demonstrating the practicality and robustness of the proposed solution.
@@ II Related Work
This section reviews key developments in fatigue assessment. Section 2.1 examines traditional and modern fatigue detection methods, highlighting their advantages and limitations, which underscore the need for a new approach that builds on strengths while addressing shortcomings. Section 2.2 explores the current data generation methods in mitigating data scarcity, along with their constraints in practical applications. Finally, Section 2.3 discusses the use of open datasets for fatigue assessment and the challenges associated with their heterogeneity and generalizability.
II-A Traditional and Diversified Fatigue Assessment
@@
Traditional fatigue assessment methods are based on subjective evaluation tools, including self-reports and questionnaires. Since the last century, numerous classic fatigue assessment scales have been widely utilized, including the Stanford Sleepiness Scale (SSS) [10], proposed by Stanford University; the Multidimensional Fatigue Inventory (MFI) [11], developed by Smets et al.; and the International Fitness Scale (IFIS) [12], which measures changes in overall vigor and affect. These scales encompass a multitude dimensions of fatigue and serve as the foundation for traditional fatigue assessment. Nevertheless, both questionnaires and self-reports are highly influenced by an individual’s psychological state and external environment, resulting in subjective and inconsistent outcomes [13]. Moreover, most of these methods [14, 15, 16] are post hoc assessments, making them unsuitable for real-time fatigue monitoring and preventive measures-both of which are essential in high-stakes environments, such as driving [17] and aviation [18].
@@
To address these limitations, researchers have explored more objective measures of human fatigue. In fact, a person’s fatigue state can be reflected in changes in physiological signals, such as heart rate (HR), heart rate variability (HRV), electrocardiogram (ECG), electroencephalogram (EEG), galvanic skin response (GSR), skin temperature (ST), blood pressure(BP), electromyography(EMG) [19, 20] and blood oxygen saturation (SpO2). Studies have shown that when drivers enter a fatigued state, their HRV metrics change significantly. Notably, a decrease is often observed in the time domain on measures such as how much the time between heartbeats naturally varies and how much consecutive heartbeats differ from each other. This suggests that the ability of the body to regulate heart activity is weakened. Similarly, HRV indicators in the frequency domain, such as the low-frequency and high-frequency ratios, often become unstable, reflecting an imbalance in the regulatory function of the autonomic nervous system [21]. Additionally, GSR signals during fatigue are usually characterized by a rise in baseline conductance level and response fluctuations, which is closely correlated to enhanced sympathetic nerve activities [22]. Through real-time monitoring and joint analysis of these specific indicators, the subtle changes in the driver’s physiological state can be captured more precisely, thus significantly improving the sensitivity and accuracy of the fatigue detection system [23]. With technological advancement, physiological signal acquisition devices are becoming smaller and more user-friendly, significantly driving research in real-time fatigue monitoring [24, 25].
@@
Among various physiological signals, EEG has gained popularity in fatigue detection due to its ability to directly measure brain activity with high accuracy. However, traditional EEG systems are expensive, require an intrusive setup and specialized personnel for operation, making it challenging for real-world deployment. As a result, the majority of EEG-based fatigue studies remain confined to controlled laboratory settings [26, 27, 28]. Although portable EEG devices have emerged for field applications, their data quality is not as good as that of laboratory-grade devices. Additionally, they are susceptible to external noise and interference, significantly impacting their stability and detection accuracy [29, 30].
@@
In response to these challenges, researchers have explored noninvasive alternatives, leading to the increase of vision-based fatigue detection [31]. These methods assess fatigue by analyzing facial features, such as eyelid closure time [32], yawning frequency [33], and body movement [34, 35] from images captured by cameras in real time. Despite their many advantages, these methods also face challenges, such as lighting variations and camera angles, which can affect image quality and ultimately detection accuracy. In addition, long-term monitoring human subjects’ facial and body movements in work environments raises privacy concerns [36] and may lead to employee resistance. Along with this trend, alternative visual signals, like pupil dilation and gaze tracking, have gained attention in the field of fatigue detection. Eyes are often considered as the most intuitive window of the mental state, which can directly reflect the immediate response to fatigue and show high accuracy in fatigue assessment. Studies have shown that pupil diameter and gaze focus are important indicators for assessing workload and fatigue status [37, 38, 39], including our own study [40] to collect human eye and physiological data for combined fatigue detection. However, while pupil dilation and gaze tracking offer promising insights, they are also sensitive to external environmental factors. Pupil dilation can be influenced by ambient lighting conditions, which may reduce its reliability in dynamic or uncontrolled environments. Similarly, gaze tracking can be affected by head movements, occlusions, and variations in camera positioning, limiting its stability and accuracy unless the camera is confined to a controlled environment, such as the one in our study [40] using Hololens.
@@
In summary, different sensor signals possess distinct advantages in fatigue detection, with some proving more effective than others. However, their effectiveness is highly contingent on applications and settings. For example, EEG is highly effective in controlled environments but impractical for real-world scenarios due to its sensitivity and complex setup requirements. Similar sensitivity issues rule out the use of pupil dilation in practical applications, and require the use of gaze tracking within a confined setup. Therefore, researchers usually select available and/or context-appropriate resources to collect signals that best align with their research objectives. Even so, we should not overlook the value of the knowledge learned from settings where these context-inappropriate sensors have been effectively applied. If a mechanism exists to transfer that knowledge, it could be used to augment fatigue detection in environments where those sensors are unavailable or impractical, thereby preserving their contribution without requiring their direct deployment.
@@ II-B Data Synthesis Methods
Data synthesis, a technology that generates artificial data that maintains the statistical or structural characteristics of real datasets, has demonstrated significant advantages in dealing with data scarcity in recent years. This is particularly relevant in physiological signal analysis, where data collection can be challenging due to various factors such as sensor availability, sensitivity to noise, and subject-related variability, leading to widespread adoption in this domain.
@@
Early developments in data synthesis(DS) primarily leveraged statistical regression techniques, where synthetic data was generated through parameterized modeling of observed data distributions. For instance, Linear regression and Bayesian regression were among the first approaches adopted to model temporal dependencies and fill in missing physiological signals by assuming probabilistic relationships among features [41]. As datasets grew in complexity and dimensionality, regularized models such as ridge regression [42] and lasso regression [43] were introduced to improve generalization. However, despite these advancements, statistical models remained fundamentally limited in their ability to capture long-term dependencies and nonlinear temporal dynamics, which are prevalent in complex physiological signals.
@@
Deep learning, on the other hand, has provided promising solutions to this challenge. Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have proven effective in capturing long-term temporal dependencies in multivariate time-series data. Several studies have demonstrated their success in a variety of applications, such as GRU-D, which incorporates decay mechanisms and masking for clinical time-series data [44]; an LSTM-based imputation model for ECG signals that outperforms classical methods [45]; and BRITS, which enables bidirectional gradient-based imputation without strong distributional assumptions [46]. In parallel, Convolutional Neural Networks (CNN) have shown great success in extracting local spatial and temporal features. Their strength has inspired hybrid models that combine CNNs with LSTMs to jointly leverage spatial and temporal dependencies for more accurate imputation and signal reconstruction [47]. Additionally, Generative Adversarial Networks (GANs) have emerged as a leading approach for producing high-fidelity synthetic signals that are statistically indistinguishable from real ones. Examples can be found in [48, 49, 50].
@@
Despite this progress, current DS techniques across both classical and deep learning paradigms typically operate under a critical limitation: they rely solely on observed data from a single modality or domain to infer or generate new data of the same type. This assumption restricts their generalizability and limits their ability to model cross-domain relationships or synthesize data from unseen modalities or sources, which is especially problematic in real-world physiological settings, where data is not only sparse and incomplete, but also highly heterogeneous across domains, such as different sensing devices, sensing modalities, and subject populations. Overcoming this constrain is essence of our work, which aims to enable more flexible and adaptive synthesis by integrating knowledge across both domains and modalities.

### v3

# Enhancing Fatigue Detection through Heterogeneous Multi-Source Data Integration and Cross-Domain Modality Imputation

@@ Abstract
Fatigue detection for human operators plays a critical role in safety critical applications such as aviation, mining, and long haul transport. While numerous studies have demonstrated the effectiveness of high-fidelity sensors in controlled laboratory environments, their performance often degrades when ported to real-world settings due to noise, lighting condition, and field of view constraints, thereby limiting their practicality. This paper formalizes a deployment-oriented setting for real-world fatigue detection, where high-quality sensors are often unavailable in practical applications. To address this challenge, we propose leveraging knowledge from heterogeneous source domains, including high-fidelity sensors that are difficult to deploy in the field but commonly used in controlled environments, to assist fatigue detection in the real-world target domain. Building on this idea, we design a heterogeneous and multi-source fatigue-detection framework that adaptively utilizes the available modalities in the target domain while exploiting diverse configurations in the source domains through cross-domain alignment and modality imputation. Our experiments, conducted using a field-deployed sensor setup and two publicly available human fatigue datasets, demonstrate the practicality, robustness, and improved generalization of our approach across subjects and domains. The proposed method achieves consistent gains over strong baselines in sensor constrained scenarios.
Note to Practitioners—The proposed work is motivated by the growing need for practical and reliable fatigue detection in industrial applications where operator fatigue poses serious safety and productivity risks. While high-performing physiological sensors like EEG provide high-fidelity signals, their use in these environments is limited by constraints such as cost, setup complexity, and sensitivity to environmental conditions. Thus, industry deployment often relies on context-appropriate sensors that may not fully capture fatigue related states. To resolve this gap, we propose a multi-source heterogeneous fatigue detection framework that enables field-deployed systems to take advantage of high-fidelity data collected from controlled environments, even if those modalities are unavailable at deployment. This is achieved through cross-domain alignment and modality imputation. Our experiments validate the practicality, robustness, and generalization of the proposed approach, showing consistent improvements in subject-independent and cross-domain settings.
@@ I Introduction
Physiological responses and behavioral patterns of individuals under fatigue provide rich information to understand the impacts of jeopardized human capacity, from which proper mitigation control can be determined. However, the complexity and vulnerability of data acquisition present several limitations. While the latest advancements of sensor technology have elevated the research of objective fatigue detection to unprecedented levels, concerns about sensor reliability and accessibility remain [Hooda2022].
@@
Unreliable sensors can result in loss of data, compromising detection accuracy, or, in the worst case, complete malfunction. To address this issue, recent efforts have explored using synthetic data generation strategies to augment incomplete or noisy datasets. For instance, generative AI [zhu2025evolutionfutureperspectivesartificial] has emerged as a promising solution. Bird et al. (2021) demonstrated the capability of GPT-2 in generating synthetic electroencephalography (EEG) and electromyography (EMG), which, when combined with real training data, improved classification accuracy [9345373]. Similarly, research in [10178849] employed generative adversarial networks (GANs) to enhance data for motor imagery classification in brain-computer interface systems. Along this direct, [zeng2021eeg] proposed Generative-DANN (GDANN), a hybrid model that integrates GANs with a Domain-Adversarial Neural Network (DANN) architecture to address cross-subject EEG variability in fatigue detection. By aligning feature distributions across subjects, GDANN improved generalizability and achieved high accuracy in cross-subject fatigue classification tasks. When generating synthetic data for fatigue detection, the process must consider inherent correlations of physiological modalities to collaboratively determine fatigue. To that end, RAINDROP [zhang2021graph] provided a promising approach by using graph neural networks to model these correlations in irregularly sampled and multivariate time-series data.
@@
Although the aforementioned approaches address the challenge of sensor unreliability through data augmentation, few efforts have tackled the equally critical issue of sensor accessibility. In many real-world scenarios, high-end sensors, such as EEG, imaging systems, and pupil dilation monitors, while reliable and accurate, remain prohibitively expensive. The extensive calibration, routine maintenance, and specialized operator training required make these sensors difficult to operate for wide, real-world deployment. Moreover, these sensors are highly sensitive to environmental factors, like noise, lighting condition, and field of view constraints. These sensitivities complicate data collection and analysis in real-world scenarios, ultimately degrading model accuracy and robustness [e19080385]. As a result, studies involving these sensors are often confined to simulated and controlled environments. For example, high resolution EEG was deployed in a simulated environment, together with heart rate and eye blink rate, to analyze mental fatigue or drowsiness during car driving [6347469]. However, such simulation based setups often fail to capture the complexity and uncertainty of actual driving conditions [8482470] , which in turn limits the generalizability of models trained in these idealized environments. Nonetheless, transferring knowledge gained from controlled settings remains valuable, provided it is adapted appropriately, offering a foundation for practical fatigue detection in real-world scenarios where direct sensor deployment is constrained. Note that this type of transfer requires a form of cross-domain adaptation that goes beyond the capabilities of existing generative methods. Current approaches, such as GANs and GPT-based models mentioined earlier, have proven effective for within-domain synthesis, where additional data is generated for modalities already present at both training and deployment (e.g., generating more EEG from partial EEG). However, they are not designed for cross-domain transfer with target-never-seen modalities, where an entire sensor modality is absent in the target domain at both training and deployment, for example, leveraging EEG-related knowledge when no EEG has ever been observed in the target system. This limitation highlights a critical gap: the need for methods that can support cross-domain sensor transfer to address real-world sensor accessibility constraints.
Motivated by these remarks, this paper proposes a novel domain-adaptive framework that enables real-world fatigue detection systems to learn from and be enhanced by controlled-environment data, even when the data originate from entirely different sensor sources. This approach offers a practical solution by facilitating cross-domain knowledge transfer, extending the utility of high-fidelity but deployment-constrained sensors to settings where they are otherwise unavailable. To this end, we present the framework with the following core contributions:
@@
  • A practical yet underexplored problem setting for real-world fatigue detection is formally defined. This definition reflects the need to operate fatigue detection systems with context-appropriate, accessible sensors in the field while leveraging knowledge transferred from differently instrumented, high-fidelity but deployment-constrained sensor sources (e.g., EEG or ECG collected in controlled lab environments). It captures the central challenges of modality heterogeneity, incompleteness, and domain shift, and lays the foundation for developing models that can generalize across mismatched sensor configurations in diverse deployment contexts.
@@
  • A heterogeneous and multi-source fatigue detection framework is proposed and validated as a solution to this problem. Grounded in a thorough theoretical analysis, we explore the conditions under which knowledge from differently instrumented source domains can be effectively leveraged to support fatigue prediction in a target domain. A domain-adaptive model is then developed to dynamically integrate available target-domain modalities while leveraging complementary information from multiple source domains. The design enables accurate fatigue prediction in real-world settings even under constrained sensor configurations. The framework and the model are validated using a practical sensor configuration and two publicly available datasets, demonstrating improved generalization and overall performance.
@@ II Related Work
This section reviews key developments in fatigue assessment. Section II.A examines existing fatigue detection methods, highlighting their advantages and limitations, which underscore the need for a new approach that builds on strengths while addressing shortcomings. Section II.B explores the current data generation methods in mitigating data scarcity, along with their constraints in practical applications. Finally, Section II.C discusses the use of open datasets for fatigue assessment and the associated challenges with sensor discrepancies and label heterogeneity.
II-A Fatigue Assessment
@@
Traditional fatigue assessment methods are based on subjective evaluation tools, including self-reports and questionnaires. Since the last century, numerous classic fatigue assessment scales have been widely utilized, including the Stanford Sleepiness Scale (SSS) [Shahid2012SSS], proposed by Stanford University; the Multidimensional Fatigue Inventory (MFI) [smets1995multidimensional], developed by Smets et al.; and the International Fitness Scale (IFIS) [Ortega2011], which measures changes in overall vigor and affect. These scales encompass a multitude dimensions of fatigue and serve as the foundation for traditional fatigue assessment. Nevertheless, both questionnaires and self-reports are highly influenced by an individual’s psychological state and external environment, resulting in subjective and inconsistent outcomes [gawron2016overview]. Moreover, most of these methods [brown2016fatigue, jap2009using, kar2010eeg] are post hoc assessments, making them unsuitable for real-time fatigue monitoring and preventive measures-both of which are essential in high-stakes environments, such as driving [app14073016] and aviation [article111].
@@
To address these limitations, researchers have explored more objective measures of human fatigue. In fact, a person’s fatigue state can be reflected in changes in physiological signals, such as heart rate (HR), heart rate variability (HRV), electrocardiogram (ECG), EEG[9464716][10488090], galvanic skin response (GSR), skin temperature (ST), blood pressure(BP), EMG [7516656, 9551478] and blood oxygen saturation (SpO2). Studies have shown that when drivers enter a fatigued state, their HRV metrics change significantly. Notably, a decrease is often observed in the time domain on measures such as how much the time between heartbeats naturally varies and how much consecutive heartbeats differ from each other. This suggests that the ability of the body to regulate heart activity is weakened. Similarly, HRV indicators in the frequency domain, such as the low-frequency and high-frequency ratios, often become unstable, reflecting an imbalance in the regulatory function of the autonomic nervous system [7473750]. Additionally, GSR signals during fatigue are usually characterized by a rise in baseline conductance level and response fluctuations, which is closely correlated to enhanced sympathetic nerve activities [Markiewicz2022]. Through real-time monitoring and joint analysis of these specific indicators, the subtle changes in the driver’s physiological state can be captured more precisely, thus significantly improving the sensitivity and accuracy of the fatigue detection system [10324398]. With technological advancement, physiological signal acquisition devices are becoming smaller and more user-friendly, significantly driving research in real-time fatigue monitoring [mohanavelu2017assessment].
@@
Among various physiological signals, EEG has gained popularity in fatigue detection due to its ability to directly measure brain activity with high accuracy. However, traditional EEG systems are expensive, require an intrusive setup and specialized personnel for operation, making it challenging for real-world deployment. As a result, the majority of EEG-based fatigue studies remain confined to controlled laboratory settings [trejo2015eeg, cheng2011mental, monteiro2019using]. Although portable EEG devices have emerged for field applications, their data quality is not as good as that of laboratory-grade devices. Additionally, they are susceptible to external noise and interference, significantly impacting their stability and detection accuracy [larocco2020systemic, shashidhar2020real].
@@
In response to these challenges, researchers have explored noninvasive alternatives, leading to the increase of vision-based fatigue detection [LIU2022105399]. These methods assess fatigue by analyzing facial features, such as eyelid closure time [10054532], yawning frequency [8808931], and body movement [10260460, 9926591] from images captured by cameras in real time. Despite their many advantages, these methods also face challenges, such as lighting variations and camera angles, which can affect image quality and ultimately detection accuracy. In addition, long-term monitoring human subjects’ facial and body movements in work environments raises privacy concerns [Shen2024] and may lead to employee resistance. Along with this trend, alternative visual signals, like pupil dilation and gaze tracking, have gained attention in the field of fatigue detection. Eyes are often considered as the most intuitive window of the mental state, which can directly reflect the immediate response to fatigue and show high accuracy in fatigue assessment. Studies have shown that pupil diameter and gaze focus are important indicators for assessing workload and fatigue status [krejtz2018eye, doi:10.1177/21695067231192895, doi:10.1076/0271-3683(200007)2111-ZFT535], including our own study [10799435] to collect human eye and physiological data for combined fatigue detection. However, while pupil dilation and gaze tracking offer promising insights, they are also sensitive to external environmental factors. Pupil dilation can be influenced by ambient lighting conditions, which may reduce its reliability in dynamic or uncontrolled environments. Similarly, gaze tracking can be affected by head movements, occlusions, and variations in camera positioning, limiting its stability and accuracy unless the camera is confined to a controlled environment, such as the one in our study [10799435] using Hololens.
@@
In summary, different sensor signals possess distinct advantages in fatigue detection, with some proving more effective than others. However, their effectiveness is highly contingent on applications and settings. For example, EEG performs well in controlled settings but can be challenging to deploy in real world contexts due to susceptibility to noise and constraints on setup and usability. Although commercial wearable EEG systems exist, practical use outside the lab often requires additional artifact handling and user training.Similar sensitivity issues rule out the use of pupil dilation in practical applications, and require the use of gaze tracking within a confined setup. Therefore, researchers usually select available and/or context-appropriate resources to collect signals that best align with their research objectives. Even so, we should not overlook the value of the knowledge learned from settings where these context-inappropriate sensors have been effectively applied. If a mechanism exists to transfer that knowledge, it could be used to augment fatigue detection in environments where those sensors are unavailable or impractical, thereby preserving their contribution without requiring their direct deployment.
@@ II-B Data Synthesis Methods
Data synthesis, a technology that generates artificial data that maintains the statistical or structural characteristics of real datasets, has demonstrated significant advantages in dealing with data scarcity in recent years. This is particularly relevant in physiological signal analysis, where data collection can be challenging due to various factors such as sensor availability, sensitivity to noise, and subject-related variability, leading to widespread adoption in this domain[NEIFAR2025103127].
@@
Early developments in data synthesis (DS) primarily leveraged statistical regression techniques, where synthetic data was generated through parameterized modeling of observed data distributions. For instance, Linear regression and Bayesian regression were among the first approaches adopted to model temporal dependencies and fill in missing physiological signals by assuming probabilistic relationships among features [kulkarni2024bayescatsivariationalbayesiandeep]. As datasets grew in complexity and dimensionality, regularized models such as ridge regression [Zahid2019mispr] and lasso regression [Deng2016] were introduced to improve generalization. However, despite these advancements, statistical models remained fundamentally limited in their ability to capture long term dependencies and nonlinear temporal dynamics, which are prevalent in complex physiological signals.
@@
Deep learning, on the other hand, has provided promising solutions to this challenge. Recurrent Neural Networks (RNNs), particularly Long Short-Term Memory (LSTM) networks, have proven effective in capturing long-term temporal dependencies in multivariate time-series data. Several studies have demonstrated their success in a variety of applications, such as GRU-D, which incorporates decay mechanisms and masking for clinical time-series data [che2016recurrent]; an LSTM-based imputation model for ECG signals that outperforms classical methods [verma2019accurate]; and BRITS, which enables bidirectional gradient-based imputation without strong distributional assumptions [cao2018brits]. In parallel, Convolutional Neural Networks (CNN) have shown great success in extracting local spatial and temporal features. Their strength has inspired hybrid models that combine CNNs with LSTMs to jointly leverage spatial and temporal dependencies for more accurate imputation and signal reconstruction [Eum2022Imputation]. Additionally, GANs have emerged as a leading approach for producing high-fidelity synthetic signals that are statistically indistinguishable from real ones. Examples can be found in [10107629, 10239214][NIA2024110129].
@@
Despite this progress, current DS techniques across both classical and deep learning paradigms typically operate under a critical limitation: they rely solely on observed data from a single modality or domain to infer or generate new data of the same type. This assumption restricts their generalizability and limits their ability to model cross domain relationships or synthesize data from unseen modalities or sources, which is especially problematic in real world physiological settings, where data is not only sparse and incomplete, but also highly heterogeneous across domains, such as different sensing devices, sensing modalities, and subject populations. Overcoming this constrain is essence of our work, which aims to enable more flexible and adaptive synthesis by integrating knowledge across both domains and modalities.